2.125 g of HNO3 will react with how many grams of copper?
1
Expert's answer
2020-05-08T14:05:27-0400

Cu + 4HNO3 → Cu(NO3)2 + 2NO2 + 2H2O;

Proportion: m(Cu)/Mr(Cu)=m(HNO3)/4*Mr(HNO3);

m(Cu)=m(HNO3)*Mr(Cu)/4*Mr(HNO3)=2.125*64/4*63=0.539 g

Answer: 0.539 g